A solution is made by adding 25mL of toluene to a beaker of benzene. If the mixture is 33% toluene, what was the volume of the benzene in the mixture?



1
Expert's answer
2021-03-12T06:29:03-0500

Benzene percentage =100-33=67%


If 33%=25mL

Then 67%=?


"=\frac {67×25}{33}=50.76mL Benzene."